# MulSMo: Multimodal Stylized Motion Generation by Bidirectional Control Flow

### v1

@@ Abstract
Generating motion sequences conforming to a target style while adhering to the given content prompts requires accommodating both the content and style. In existing methods, the information usually only flows from style to content, which may cause conflict between the style and content, harming the integration. Differently, in this work we build a bidirectional control flow between the style and the content, also adjusting the style towards the content, in which case the style-content collision is alleviated and the dynamics of the style is better preserved in the integration. Moreover, we extend the stylized motion generation from one modality, i.e. the style motion, to multiple modalities including texts and images through contrastive learning, leading to flexible style control on the motion generation. Extensive experiments demonstrate that our method significantly outperforms previous methods across different datasets, while also enabling multimodal signals control. The code of our method will be made publicly available.
@@ 1 Introduction
Figure 2: Overview of MulSMo. Our approach can take in various modalities as style signals, such as style motion sequences, texts, and images. We generate stylized human motions by combining content text $\mathbf{c}$ with style signals $\mathbf{s}$. In the encoder part of the latent denoiser, we establish a bidirectional control flow between the generation network and the style network, utilizing respective zero linear for fusion. (Note that for simplicity we plot the style-to-content zero linear and content-to-style zero linear into one block in the figure.) This denoising step is repeated $T$ times to obtain the noise-free motion latent ${\bm{z}}_{0}$, which is decoded to stylized motion by the decoder.
@@
To address these two problems, in this paper, we propose a Multimodal Stylized Motion generation model (MulSMo), where a bidirectional control mechanism is designed to impose the style onto content motion, and the contrastive learning is leveraged to enable multi-modality style control, as shown in Figure 1. Specifically, we establish a bidirectional control flow between the style network and generation network, such that the control not only flows from the style to content, but also flows from the content to style, as displayed in Figure 2. This enables the style to also adjust towards the content, bridging the gap between two spaces and relieving the conflict, especially the divergence in dynamics. To enable both local and global bidirectional control, we equip the bidirectional flow modules in each encoder block so that local features can be controlled in the shallow layers and more global features is controlled while going deeper. To enable multimodal signals for style control, we propose to align the style embeddings across style motion sequences, text descriptions, and images within a contrastive learning manner. In this way, the style text descriptions or style images can play the same role as style motion sequences for style control during generation. Extensive experiments are performed with either style motion, texts, or images as style input, and demonstrate that our method obtains better performance compared to previous methods.
@@ 3.1 Overview
In this section, we present our proposed MulSMo, which generates stylized motion sequences by taking text descriptions and multi-modal style signals as input. An overview of our model is depicted in Figure 2. Aligning with the settings in MLD [6], we also place the diffusion process in the motion latent space. In our method, $\epsilon_{\theta}$ denotes the latent denoiser and $\{{\bm{z}}_{t}\}_{t=0}^{T}$ represents the noised latent sequences, where ${\bm{z}}_{T}$ is a Gaussian noise. Our purpose is to denoise from ${\bm{z}}_{T}$ to a clean stylized latent motion sequence. Given a content prompt $\mathbf{c}$ and a style signal $\mathbf{s}$, we define the denoising process at step $t\in(0,T]$ as $\epsilon_{t}=\epsilon_{\theta}({\bm{z}}_{t},t,\mathbf{c},\mathbf{s})$, where $\epsilon_{\theta}$ denotes the predicted noise at timestep $t$. After denoising for $T$ timesteps, we obtain a clean latent ${\bm{z}}_{0}$, which is then decoded by a pretrained decoder into a stylized motion sequence $\hat{{\bm{x}}}_{0}\in\mathbb{R}^{L\times D}$, where $L$ denotes the frame length of motion and $D$ is the dimension of human motion representations. Following [46], we adopt the same motion representations as detailed in HumanML3D [11], with $D=263$. As depicted in Figure 2, the text description $\mathbf{c}$ is used to control the content domain of the generated motion, while the style domain is controlled by style signals, including reference motion sequences, texts, or images.
@@ 3.3 Multimodal Style Control
As shown in Figure 4, we input the style label into CLIP to obtain text embeddings $t$. After passing through an adaptor, we align the text embeddings $t$ with the style embeddings $s$ and optimize them using the InfoNCE loss [14]:
$\mathcal{L}_{\text{InfoNCE}}=-\log\frac{\exp(\text{sim}(t_{i},s_{j})/\tau_{t})}{\sum_{k=1}^{N}\exp(\text{sim}(t_{i},s_{k})/\tau_{t})},$ (3)
where $t_{i}$ and $s_{j}$ are positive samples, while $t_{i}$ and $s_{k}$ are negative samples. sim($\cdot$) means cosine similarity and $\tau$ denotes the temperature hyperparameter. During the inference stage, we can use a single label word as a textual control signal for stylized motion generation. Since CLIP has already aligned language and vision, we can also input images as control signals, which are processed through the adaptor to obtain style-aware image embeddings as style control.
@@ 3.4 Stylized Motion Generation Guidance
Ours 77.042 1.527 4.292 0.613 9.303 0.118
SMooDi [46] 81.298 3.066 5.011 0.439 9.433 0.155
Ours 86.455 3.028 4.872 0.498 9.514 0.153
@@
Adaptor+MLD (Ours) 52.100 0.754 3.875 0.674 8.875 0.104
@@ 3.5 Learning Scheme
We freeze the parameters of MLD and only train the style network on the 100STYLE dataset using the following loss function:
@@
Overall, the training loss function of our framework is defined as follows:
@@ 4.1 Experiment Setup
Implementation Details.
We utilize MLD as the pretrained generation network and train the style network from scratch. The style motion encoder in Figure 4 consists of a single transformer encoder, and the adaptor is a simple MLP. During the training of the style control network, the style network composed of four Transformer Encoder blocks is optimized while the parameters of the MLD are frozen. The framework is trained with the AdamW optimizer [23] and the learning rate is 1e-5. More details can be found in the supplementary.
Datasets.
Evaluation Metrics.
The metrics are designed to assess three key dimensions: Content Preservation, Style Reflection, and Realism. For evaluating content preservation, we follow [6] and use metrics including motion-retrieval precision (R-precision), Multi-modal Distance (MM Dist), Diversity, and Frechet Inception Distance (FID). For realism, we address common foot skating issues by integrating the Foot Skating Ratio metric proposed by [20]. For style reflection, we follow SMooDi [46] to use Style Recognition Accuracy (SRA) [17]. SRA and FID are two metrics that involve a trade-off. A higher SRA indicates more pronounced stylization but leads to a relatively worse FID. During the evaluation, a content text from the HumanML3D dataset and a motion style sequence from the 100STYLE dataset, are randomly selected to generate the stylized motion. A pretrained style classifier then computes the SRA for the generated motion, following previous methods [46].
@@
Baselines.
We focus on three primary tasks: motion-guided stylized motion generation, motion style transfer, and text/image-guided stylized motion generation. For the motion-guided stylized motion generation and motion style transfer tasks, we utilize the evaluation metrics established in SMooDi [46], incorporating MLD as the text-to-motion model; the methods for motion style transfer are consistent with those employed in the motion style transfer task. For the text/image-guided stylized motion generation task, we extract features from the texts and images using CLIP and input these features into our trained adaptor as style embeddings for style control. Unlike ChatGPT+MLD, which essentially functions as a straightforward text-to-motion model without embodying the control aspect, our approach allows for style control using a simple label word, enabling more nuanced manipulation of the generated motion styles.
@@ 4.2 Evaluation
Motion-guided stylized motion generation.
For the stylized text-to-motion generation task with motions as style, we present a comparative analysis of our approach against four baseline methods, as summarized in Table 1. When we input style motion as the control signal, our model outperforms existing methods on both SRA and FID metrics. SRA and FID are two trade-off metrics; however, compared to SMooDi [46], our approach achieves an FID that is lower by 5.1% and 1.3% while maintaining an SRA that is 6.4% higher. As illustrated in Figure 6, SMooDi exhibits conflicts between style and content, resulting in the stylized motion failing to align with the content prompts. Furthermore, the style motion tends to lose some dynamic qualities, leading to inaccuracies in classifying stylized actions for similar poses, such as ‘Flapping’ and ‘Aeroplane’.
Text/image-guided stylized motion generation.
When the control signal comes to texts, ChatGPT+MLD utilizes ChatGPT to merge style labels from 100STYLE with texts from HumanML3D into a sentence. For example, ‘a person walks’ + ‘old’ = ‘an elderly person walks’. Then this merged sentence is fed to MLD. The performance of ChatGPT+MLD is limited, achieving merely 4.819% in the SRA metric. This suggests that the MLD framework, despite incorporating style descriptors, falls short in facilitating stylized generation from text inputs alone. In contrast, our Adaptor+MLD accepts one label word as signal control, and fully leverages the control mechanism of the MulSMo, improving the SRA performance from 4.819% to 52.1% with only a small degradation on the FID. The visualization of the generated stylized motions from texts or images is displayed in Figure 5, from where we can see the convenient multimodal control.
Motion style transfer.
For the motion style transfer task, we follow the settings of SMooDi and evaluate only three metrics: FID, SRA, and Foot skating ratio. We utilize the HumanML3D dataset [11] as the source of motion content and extract motion styles from the 100STYLE dataset. As shown in Table 2, our algorithm achieves a significant improvement of 7.8% on the SRA metric, along with advancements of 1.0% and 8.4% on the FID and Foot skating ratio metrics, respectively. Furthermore, we conduct experiments on the Xia dataset [40], which was not seen during the training phase of both our model and the baseline model. Given the presence of motion content labels in the Xia dataset, we report the Content Recognition Accuracy (CRA). The results are presented in Table 3.
Figure 7: User study on the stylized motion generation task.
@@ 4.3 Ablation Study
In Figure 3 (a), we align our approach with SMooDi, introducing a unidirectional control signal between the encoder of the generation network and the encoder of the style network, with results corresponding to the first row in Table 4. In Figure 3 (b), we implement a bidirectional control flow between these two encoders, where the style signal influences the content, and at the same time, the content signal provides feedback to the style. We choose this as the final framework. In Figures 3 (c) and 3 (d), bidirectional signals are also incorporated between the middle block and the decoder, respectively. These two frameworks lead to a degradation in the final results. We attribute this to the control signals from the style network disrupting the parameters of the frozen middle block and decoder of the generation network, thereby affecting the recovery of fine details.
4.4 User Study
In order to address the inherently subjective nature of stylized motion, we conduct user studies employing pairwise comparisons to systematically evaluate our method in the task of stylized motion generation. A total of 41 human subjects were recruited to participate in this study. During each trial, each participant will be provided with two video segments, each approximately 9 seconds in length, generated by our method and SMooDi, respectively. They are tasked with selecting their preferred video, taking into account three key dimensions: Realism, Style Reflection, and Content Preservation. As illustrated in Figure 7, our method garnered significantly higher user preference across all three dimensions when compared to the SMooDi. Details regarding the study are provided in the supplementary materials.
@@ 5 Conclusion
This paper proposes a multimodal stylized motion generation framework that innovatively leverages style motion, text, or image as style control to produce dynamic motion sequences. Different from previous methods where the control information only moves from style to content, a bidirectional control flow between the style network and the generation network is also introduced to resolve the conflict between the two spaces for better stylized motion generation.

### v2

@@ Abstract
Generating motion sequences conforming to a target style while adhering to the given content prompts requires accommodating both the content and style. In existing methods, the information usually only flows from style to content, which may cause conflict between the style and content, harming the integration. Differently, in this work we build a bidirectional control flow between the style and the content, also adjusting the style towards the content, in which case the style-content collision is alleviated and the dynamics of the style is better preserved in the integration. Moreover, we extend the stylized motion generation from one modality, i.e. the style motion, to multiple modalities including texts and images through contrastive learning, leading to flexible style control on the motion generation. To further boost the performance, we advance the motion diffusion to motion-aligned temporal latent diffusion by developing a novel motion VAE. Extensive experiments demonstrate that our method significantly outperforms previous methods across different datasets, while also enabling multimodal signals control. The code of our method will be made publicly available.
@@ 1 Introduction
Figure 2: Overview of MulSMo. Our approach can take in various modalities as style signals, such as style motion sequences, texts, and images. We generate stylized human motions by combining content text $\mathbf{c}$ with style signals $\mathbf{s}$. In the encoder part of the latent denoiser, we establish a bidirectional control flow between the generation network and the style network, utilizing respective zero linear for fusion. (Note that for simplicity we plot the style-to-content zero linear and content-to-style zero linear into one block in the figure.) This denoising step is repeated $T$ times to obtain the noise-free motion latent ${\bm{z}}_{0}$, which is decoded to stylized motion by the latent decoder.
@@
To address these two problems, in this paper, we propose a Multimodal Stylized Motion generation model (MulSMo), where a bidirectional control mechanism is designed to impose the style onto content motion, and the contrastive learning is leveraged to enable multi-modality style control, as shown in Figure 1. Specifically, we establish a bidirectional control flow between the style network and generation network, such that the control not only flows from the style to content, but also flows from the content to style, as displayed in Figure 2. This enables the style to also adjust towards the content, bridging the gap between two spaces and relieving the conflict, especially the divergence in dynamics. To enable both local and global bidirectional control, we equip the bidirectional flow modules in each encoder block so that local features can be controlled in the shallow layers and more global features is controlled while going deeper. To enable multimodal signals for style control, we propose to align the style embeddings across style motion sequences, text descriptions, and images within a contrastive learning manner. In this way, the style text descriptions or style images can play the same role as style motion sequences for style control during generation.
To further boost the performance, we also advance the latent diffusion to motion-aligned temporal latent diffusion (MaTLD) by designing a Motion-aligned Temporal VAE, which employs a temporal gating mechanism to enhance sequential pattern preservation and aligns the motion embedding towards motion priors. Extensive experiments are performed with either style motion, texts, or images as style input, and demonstrate that our method obtains better performance compared to previous methods.
@@
  • We advance the diffusion generation to motion-aligned temporal latent diffusion, which leads to a substantial enhancement in the motion generation.
@@ 3.1 Overview
In this section, we present our proposed MulSMo, which generates stylized motion sequences by taking text descriptions and multi-modal style signals as input. An overview of our model is depicted in Figure 2. Aligning with MLD [6], we propose MaTLD to operate the diffusion process in the motion latent space with motion-aligned temporal VAE, which is depicted in Figure 3. In our method, $\epsilon_{\theta}$ denotes the latent denoiser and $\{{\bm{z}}_{t}\}_{t=0}^{T}$ represents the noised latent sequences, where ${\bm{z}}_{T}$ is a Gaussian noise. Our purpose is to denoise from ${\bm{z}}_{T}$ to a clean stylized latent motion sequence. Given a content prompt $\mathbf{c}$ and a style signal $\mathbf{s}$, we define the denoising process at step $t\in(0,T]$ as $\epsilon_{t}=\epsilon_{\theta}({\bm{z}}_{t},t,\mathbf{c},\mathbf{s})$, where $\epsilon_{\theta}$ denotes the predicted noise at timestep $t$. After denoising for $T$ timesteps, we obtain a clean latent ${\bm{z}}_{0}$, which is then decoded by a pretrained decoder into a stylized motion sequence $\hat{{\bm{x}}}_{0}\in\mathbb{R}^{L\times D}$, where $L$ denotes the frame length of motion and $D$ is the dimension of human motion representations. Following [50], we adopt the same motion representations as detailed in HumanML3D [11], with $D=263$. As depicted in Figure 2, the text description $\mathbf{c}$ is used to control the content domain of the generated motion, while the style domain is controlled by style signals, including reference motion sequences, texts, or images.
@@
3.2 Motion-aligned Temporal VAE
In the context of the MLD-based Variational Autoencoder (VAE) [6], the encoder performs resampling on motion sequences of varying lengths to produce a single latent motion token, upon which a diffusion process is conducted. However, this compression along the temporal dimension results in the loss of certain temporal information within the latent token. To mitigate this issue, we employ a gating mechanism to attenuate noise introduced into the attention layers depicted in Figure 3, allowing the model to concentrate more effectively on interactions occurring at critical time steps. Specifically, for each layer of the transformer, prior to the self-attention operation, we apply an MLP to the input features, followed by a softmax operation to compute weights. These weights are then multiplied with the input features to reduce the temporal information loss associated with the resampled single motion token while simultaneously serving to pre-filter noise (e.g., decreasing the weights of frames corresponding to stagnant motion).
Due to the training strategy of the VAE, the latent representations tend to quickly deviate from their initial states in pursuit of reconstruction optimization, which may limit their expressiveness. To address this issue, we align the motion latent tokens with the motion transformer of the LaMP [24], aiming to enhance the expressive capacity of the latent representations while simultaneously regularizing the high-dimensional latent space. Furthermore, diffusion models generate latent vectors through denoising, and their outputs may carry residual noise and a certain degree of stochasticity, which does not align with the clean latent representations expected by the decoder of VAE. Therefore, we posit that this alignment process can also be regarded as an effective means of enhancing the robustness of the decoder. Our proposed motion-aligned temporal VAE is optimized with KL loss, reconstruction loss [23], and InfoNCE loss [14] for alignment:
$\mathcal{L}=\mathcal{L}_{\text{KL}}+\mathcal{L}_{\text{recon}}+\mathcal{L}_{\text{InfoNCE}}$ (1a)
$\mathcal{L}_{\text{InfoNCE}}=-\log\frac{\exp(\text{sim}(t_{i},s_{j})/\tau_{t})}{\sum_{k=1}^{N}\exp(\text{sim}(t_{i},s_{k})/\tau_{t})},$ (1b)
where $t_{i}$ and $s_{j}$ are positive samples, while $t_{i}$ and $s_{k}$ are negative samples. sim($\cdot$) means cosine similarity and $\tau$ denotes the temperature hyperparameter.
@@ 3.4 Multimodal Style Control
As shown in Figure 5, we input the style label into CLIP to obtain text embeddings $t$. After passing through an adaptor, we align the text embeddings $t$ with the style embeddings $s$ and optimize them using the InfoNCE loss. During the inference stage, we can use a single label word as a textual control signal for stylized motion generation. Since CLIP has already aligned language and vision, we can also input images as control signals, which are processed through the adaptor to obtain style-aware image embeddings as style control.
@@ 3.5 Stylized Motion Generation Guidance
BiFlow + MLD 77.042 1.527 4.292 0.613 9.303 0.118
BiFlow + MaTLD (Ours) 78.029 1.511 3.823 0.676 8.939 0.116
Motion SMooDi [50] 81.298 3.066 5.011 0.439 9.433 0.155
BiFlow + MLD 86.455 3.028 4.872 0.498 9.514 0.153
BiFlow + MaTLD (Ours) 87.324 2.977 4.636 0.543 8.115 0.149
@@
Adaptor + MLD 52.100 0.754 3.875 0.674 8.875 0.104
Adaptor + MaTLD (Ours) 52.307 0.749 3.587 0.710 8.687 0.114
@@ 3.6 Learning Scheme
We freeze the parameters of MaTLD and only train the style network on the 100STYLE dataset using the following loss function:
@@
Overall, the training loss function is defined as follows:
@@ 4.1 Experiment Setup
We utilize MaTLD (pretrained with LaMP [24]) as the generation network and train the style network from scratch. The style motion encoder in Figure 5 consists of a single transformer encoder, and the adaptor is a simple MLP. During the training of the style control network, the style network composed of four Transformer Encoder blocks is optimized while the parameters of the MaTLD are frozen. The framework is trained with the AdamW optimizer [25] and the learning rate is 1e-5. We evaluate motion generation, stylized motion generation, and motion style transfer on HumanML3D [11], KIT-ML [29], 100STYLE [26], and Xia [44] datasets. More details about the datasets, metrics, and baselines can be found in the supplementary.
@@
Top 1 Top 2 Top 3
Dataset Ground Truth $0.511^{\pm.003}$ $0.703^{\pm.003}$ $0.797^{\pm.002}$ $0.002^{\pm.00}$ $2.974^{\pm.008}$ $9.503^{\pm.065}$
HumanML3D MLD [6] $0.481^{\pm.003}$ $0.673^{\pm.003}$ $0.772^{\pm.002}$ $0.473^{\pm.013}$ $3.196^{\pm.010}$ $9.724^{\pm.082}$
MaTLD w. LaMP $0.542^{\pm.003}$ $0.733^{\pm.002}$ $0.828^{\pm.001}$ $0.152^{\pm.010}$ $2.878^{\pm.006}$ $9.703^{\pm.078}$
Dataset Ground Truth $0.424^{\pm.005}$ $0.649^{\pm.006}$ $0.779^{\pm.006}$ $0.031^{\pm.004}$ $2.788^{\pm.012}$ $11.080^{\pm.097}$
KIT-ML MLD [6] $0.390^{\pm.008}$ $0.609^{\pm.008}$ $0.734^{\pm.007}$ $0.404^{\pm.027}$ $3.204^{\pm.027}$ $10.80^{\pm.117}$
MaTLD w. LaMP $0.454^{\pm.007}$ $0.653^{\pm.006}$ $0.771^{\pm.006}$ $0.294^{\pm.015}$ $2.909^{\pm.024}$ $10.623^{\pm.107}$
Table 4: The quantitative results of our MaTLD with evaluator following previous methods on the HumanML3D dataset and the KIT-ML dataset, compared with MLD.
[…]
@@ 4.2 Evaluation
Motion-guided stylized motion generation. For the stylized text-to-motion generation task with motions as style, we present a comparative analysis of our approach against four baseline methods, as summarized in Table 1. When we input style motion as the control signal, our model outperforms existing methods on both SRA and FID metrics. SRA and FID are two trade-off metrics; however, compared to SMooDi [50], our approach achieves an FID that is lower by 6.1% while maintaining an SRA that is 7.7% higher. As illustrated in Figure 7, SMooDi exhibits conflicts between style and content, resulting in the stylized motion failing to align with the content prompts. Furthermore, the style motion tends to lose some dynamic qualities, leading to inaccuracies in classifying stylized actions for similar poses, such as ‘Flapping’ and ‘Aeroplane’.
Text/image-guided stylized motion generation. When the control signal comes to texts, ChatGPT+MLD utilizes ChatGPT to merge style labels from 100STYLE with texts from HumanML3D into a sentence. For example, ‘a person walks’ + ‘old’ = ‘an elderly person walks’. Then this merged sentence is fed to MLD. The performance of ChatGPT+MLD is limited, achieving merely 4.819% in the SRA metric. This suggests that the MLD framework, despite incorporating style descriptors, falls short in facilitating stylized generation from text inputs alone. In contrast, our Adaptor+MLD/MaTLD accepts one label word as signal control, and fully leverages the control mechanism of the MulSMo, improving the SRA performance from 4.819% to 52.1% and 52.307% with only a small degradation on the FID. The visualization of the generated stylized motions from texts or images is displayed in Figure 6, from where we can see the convenient multimodal control.
Motion style transfer. For the motion style transfer task, we follow the settings of SMooDi and evaluate only three metrics: FID, SRA, and Foot skating ratio. We utilize the HumanML3D dataset [11] as the source of motion content and extract motion styles from the 100STYLE dataset. As shown in Table 2, our algorithm achieves a significant improvement of 9.4% on the SRA metric, along with advancements of 5.8% and 8.4% on the FID and Foot skating ratio metrics, respectively. Furthermore, we conduct experiments on the Xia dataset [44], which was not seen during the training phase of both our model and the baseline model. Given the presence of motion content labels in the Xia dataset, we report the Content Recognition Accuracy (CRA). The results are presented in Table 3.
Text-to-Motion Generation. In Table 4.1, we report the text-to-motion results using CLIP and LaMP as text encoders, comparing them with MLD. We employ our proposed motion-aligned temporal VAE encoding to obtain the motion latent representations, and the framework of the diffusion process is consistent with MLD. As shown in the table, there is a significant improvement compared to MLD, such as an improvement of 6.9%, 5.3%, and 4.7% in R-Precision Top 1, 2, 3 and improves FID by 28.9% on HumanML3D, indicating that our enhancements to the VAE are both reasonable and effective.
@@ 4.3 Ablation Study
In Figure 4 (a), we align our approach with SMooDi, introducing a unidirectional control signal between the encoder of the generation network and the encoder of the style network, with results corresponding to the first row in Table 5. In Figure 4 (b), we implement a bidirectional control flow between these two encoders, where the style signal influences the content, and at the same time, the content signal provides feedback to the style. We choose this as the final framework. In Figures 4 (c) and 4 (d), bidirectional signals are also incorporated between the middle block and the decoder, respectively. These two frameworks lead to a degradation in the final results. We attribute this to the control signals from the style network disrupting the parameters of the frozen middle block and decoder of the generation network, thereby affecting the recovery of fine details.
@@ 5 Conclusion
This paper proposes a motion-aligned temporal VAE and a multimodal stylized motion generation framework that innovatively leverages style motion, text, or image as style control to produce dynamic motion sequences. Different from previous methods where the control information only moves from style to content, a bidirectional control flow between the style network and the generation network is also introduced to resolve the conflict between the two spaces for better stylized motion generation.